OpenStudy (anonymous):

The sum of a number and 6 times its reciprocal is 7. Find the number. If the problem has two solutions, type your answers in the box below with NO spaces and separating them with a comma. Example: 2,7

OpenStudy (lgbasallote):

hint \[\large x + 6(\frac{1}{x}) = 7\]

OpenStudy (anonymous):

In symbols, your first sentence is this, where \(x\) is your "number".\[\Huge x+6\cdot \frac{1}{x}=7\]Now solve for \(x\) (Hint: multiply both sides by \(x\) first and then put all the terms on one side.)

OpenStudy (anonymous):

x=1,6? :o

OpenStudy (anonymous):

Good job! That is correct. Don't forget that to check your answers you can always substitute back into the equation to see if the solutions hold true.

OpenStudy (anonymous):

Thanks so much c: I have a few more on the way 'xD <3